# Towards a Scalable Reference-Free Evaluation of Generative Models

### v1

@@ 6 Numerical Results
We evaluated the FKEA method on several image, text, and video datasets to assess its performance in quantifying diversity in different domains. In the experiments, we computed the empirical covariance matrix of $2r$-dimensional Fourier features with a Gaussian kernel with bandwidth parameter $\sigma$ tuned for each dataset, and then applied FKEA approximation for the $\mathrm{VENDI}_{1}$, $\mathrm{VENDI}_{1.5}$, and the $\mathrm{RKE}$ (same as $\mathrm{VENDI}_{2}$) scores. Experiments were conducted on 8 RTX3090 GPUs. We interpreted the modes identified by FKEA entropy-based diversity evaluation using the eigenvectors of the proxy covariance matrix as discussed in Remark 2. For each eigenvector, we presented the training data with maximum eigenfunction values corresponding to the eigenvector.
6.1 Experimental Results on Image Data
@@
To investigate the FKEA method’s diversity evaluation in settings where we know the ground-truth clusters and their quantity, we simulated an experiment on the colored MNIST [46] data with the images of 10 colored digits as shown in Figure 2. We evaluated the FKEA approximations of the diversity scores while including samples from $t$ digits for $t\in\{1,\ldots,10\}$. The plots in Figure 2 indicate the increasing trend of the scores and FKEA’s tight approximations of the scores. Also, we show the top-25 training samples with the highest scores according to the top-10 FKEA eigenvectors, showing the method captured the ground-truth clusters.
@@
We conducted an experiment on the ImageNet data to monitor the scores’ behavior evaluated for 50k samples from an increasing number of ImageNet labels. Figure 3 shows the increasing trend of the scores as well as the top-9 samples representing the top-4 identified clusters used for the entropy calculation. Figures 3, 5 and 4 illustrate the top FKEA modes with corresponding parameters. The modes represent a specific characteristic common between the top images. Table 1 evaluates the diversity scores of various ImageNet GAN models using the FKEA method applied to VENDI-1 and RKE, with potential extension to the entire VENDI family. The comparison includes baseline diversity metrics such as Inception Score [12], FID [7], Improved Precision/Recall [10], and Density/Coverage [11]. Also, Figure 7 presents the FKEA approximated entropy scores with different truncation factors in StyleGAN3 [47] on 30k generated data for each truncation factor, showing the increasing diversity scores with the truncation factor.
We defer discussing the results on generated synthetic images to the Appendix.
@@
Figure 3: RFF-based identified clusters used in FKEA Evaluation in ImageNet[6] dataset with DinoV2 embedding, Fourier feature dimension $2r=16k$ and Gaussian Kernel bandwidth $\sigma=25$. The graphs indicate increase in FKEA diversity metrics with increasing number of labels per 50k samples.
Figure 4: RFF-based identified clusters used in FKEA Evaluation in FFHQ[48] dataset with DinoV2 embedding, Fourier feature dimension $2r=16k$ and Gaussian Kernel bandwidth $\sigma=20$.
Figure 5: RFF-based identified clusters used in FKEA Evaluation in AFHQ[49] dataset with DinoV2 embedding, Fourier feature dimension $2r=16k$ and Gaussian Kernel bandwidth $\sigma=20$.
Figure 6: RFF-based identified clusters used in FKEA Evaluation in MS COCO [50] dataset with DinoV2 embedding, Fourier feature dimension $2r=16k$ and Gaussian Kernel bandwidth $\sigma=22$.
@@
Dataset (100k) - - - - - - 9176.9 996.7
ADM [51] 542.6 11.12 0.78 0.79 0.88 0.89 8360.3 633.4
ADMG [51] 659.3 5.63 0.87 0.84 0.80 0.85 8524.2 811.5
ADMG-ADMU [51] 701.6 4.78 0.90 0.73 1.20 0.96 8577.6 839.8
BigGAN [52] 696.4 7.91 0.81 0.44 0.99 0.57 7120.5 492.4
DiT-XL-2 [53] 743.2 3.56 0.92 0.84 1.16 0.97 8626.5 855.8
GigaGAN [54] 678.8 4.29 0.89 0.74 0.74 0.70 8432.5 671.6
LDM [55] 734.4 4.75 0.93 0.76 1.04 0.93 8573.7 811.9
Mask-GIT [56] 717.4 5.66 0.91 0.72 1.01 0.82 8557.4 759.5
RQ-Transformer [57] 558.3 9.57 0.80 0.76 0.77 0.59 8078.4 512.1
StyleGAN-XL[58] 675.4 4.34 0.89 0.74 1.18 0.96 8171.9 703.5
6.2 Experimental Results on Text and Video Data
To perform experiments on the text data with known clustering ground-truth, we generated 500,000 paragraphs using GPT-3.5 [59] about 100 randomly selected countries (5k samples per country). In the experiments, the text embedding used was text-embedding-3-large [60, 61, 59]. We evaluated the diversity scores over data subsets of size 50k with different numbers of mentioned countries. Figure 8 shows the growing trend of the diversity scores when including more countries. The figure also shows the countries mentioned in the top-6 modes provided by FKEA-based principal eigenvectors, which shows the RFF-based clustering of countries correlates with their continent and geographical location. We discuss the numerical results on non-synthetic text datasets, CNN/Dailymail [62][63], CMU Movie Corpus [64], in the Appendix.
In Table 2 we extend FKEA to non-synthetic Wikipedia dataset. To visualise the results, we used YAKE [65] algorithm to extract and present the identified unigram and bigram keywords. Identified mode 1 correlates most with historical figures/events/places. Mode 2 clusters smaller villages and rural regions together. Mode 3 is exclusively about people in sports, such as athletes and referees. Mode 4 visualises various music bands and albums. Lastly, mode 5 presents the articles about sports events, such as football leagues.
Mode #1 Mode #2 Mode #3 Mode #4 Mode #5
Clay Huffman Zasun Dmitri Nezhelev EX-Girl 2020 HFX Wanderers FC…
Keywords
United States village World Cup Records albums League
World War Osh Region Summer Olympics Singles Chart League Soccer
For video data experiments, we considered two standard video datasets, UCF101 [66] and Kinetics-400 [67]. Following the video evaluation literature [68, 69], we used the I3D pre-trained model [70] as embedding, which maps a video sample to a 1024-dimensional vector. As shown in Figure 9, increasing the number of video classes of test samples led to an increase in the FEKA approximated diversity metrics. Also, while the samples identified for the first identified cluster look broad, the next modes seemed more specific. We discuss the results of the Kinetics dataset in the Appendix.
@@ 6.2 Experimental Results on Text and Video Data
7 Limitations
@@ A.1 Proof of Theorem 1
[…]
@@
[…]
@@
[…]
@@
Assume $\mathrm{Tr}\rho^{\alpha}$ (resp., $-\mathrm{Tr}\rho\log\rho$) has circuit size $s(d)$. By Lemma 1 and Lemma 3, $\nabla\mathrm{Tr}\rho^{\alpha}=\alpha\rho^{\alpha-1}$ (resp., $-\nabla\mathrm{Tr}\rho\log\rho=\log\rho-1/\ln 2$) has circuit size $O(s(d))$. For every symmetric matrix $X$ and sufficiently small $t$, the matrix $\rho=I+tX$ is positive semi-definite. By Lemma 2 the $R^{d^{2}}$-valued function $\partial^{2}\rho/\partial t^{2}$ has circuit size $O(s^{d})$. The value of this function at $t=0$ is $\alpha(\alpha-1)(\alpha-2)X^{2}$ (resp., $X^{2}$), namely the square of the input matrix $X$ up to constant. Finally, computing the product $AB$ reduces to squaring the symmetric matrix
@@ A.2 Proof of Theorem 2
=$\stackrel{{\scriptstyle(a)}}{{=}}\,\int\widehat{\kappa_{\sigma}}(\bm{\omega})\exp\bigl{(}i\bm{\omega}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\mathrm{d}\bm{\omega}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\,\int\widehat{\kappa_{\sigma}}(\bm{\omega})\cos\bigl{(}\bm{\omega}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\mathrm{d}\bm{\omega}$
$\stackrel{{\scriptstyle}}{{=}}\,\mathbb{E}_{\bm{\omega}\sim\widehat{\kappa}}\Bigl{[}\cos\bigl{(}\bm{\omega}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\Bigr{]}$
$=\,\mathbb{E}_{\bm{\omega}\sim\widehat{\kappa}}\Bigl{[}\cos\bigl{(}\bm{\omega}^{\top}\mathbf{x})\cos\bigl{(}\bm{\omega}^{\top}\mathbf{x}^{\prime})+\sin\bigl{(}\bm{\omega}^{\top}\mathbf{x})\sin\bigl{(}\bm{\omega}^{\top}\mathbf{x}^{\prime})\Bigr{]}$
@@
Therefore, since $\bigl{|}\cos\bigl{(}\bm{\omega}^{\top}\mathbf{y}\bigr{)}\bigr{|}\leq 1$ for all $\bm{\omega}$ and $\mathbf{y}$, one can apply Hoeffding’s inequality to show for independently drawn $\bm{\omega}_{1},\ldots,\bm{\omega}_{r}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\widehat{\kappa}$ the following probably correct bound holds:
$\mathbb{P}\biggl{(}\Bigl{|}\frac{1}{r}\sum_{i=1}^{r}\cos\bigl{(}\bm{\omega}_{i}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}-\mathbb{E}_{\bm{\omega}\sim\widehat{\kappa}}\Bigl{[}\cos\bigl{(}\bm{\omega}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\Bigr{]}\Bigr{|}\geq\epsilon\biggr{)}\leq 2\exp\Bigl{(}-\frac{r\epsilon^{2}}{2}\Bigr{)}$
Therefore, as the identity $\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)$ reveals $\frac{1}{r}\sum_{i=1}^{r}\cos\bigl{(}\bm{\omega}_{i}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\,=\,\widetilde{\phi}_{r}(\mathbf{x})^{\top}\widetilde{\phi}_{r}(\mathbf{x}^{\prime})$, the above bound can be rewritten as
@@
$\mathbb{P}\biggl{(}\max_{1\leq i,j\leq n}\Bigl{(}\widetilde{\phi}_{r}(\mathbf{x}_{i})^{\top}\widetilde{\phi}_{r}(\mathbf{x}_{j})-k_{\text{\rm Gaussian}(\sigma^{2})}(\mathbf{x}_{i},\mathbf{x}_{j})\Bigr{)}^{2}\geq\epsilon^{2}\biggr{)}\leq 2{n\choose 2}\exp\Bigl{(}-\frac{r\epsilon^{2}}{2}\Bigr{)}.$
@@
$\mathbb{P}\Bigl{(}\bigl{\|}\frac{1}{n}\widetilde{K}-\frac{1}{n}K\bigr{\|}^{2}_{F}\,\geq\,n^{2}\frac{\epsilon^{2}}{n^{2}}\Bigr{)}\;\leq\;{n\choose 2}\exp\Bigl{(}-\frac{r\epsilon^{2}}{2}\Bigr{)}.$
@@
Appendix B Additional Numerical Results
B.1 Synthetic Image Dataset Modes
In addition to running clustering on real image datasets, we also applied FKEA with varying Gaussian Kernel bandwidth parameter $\sigma$ to generated synthetic datasets. The results are presented for LDM, VDVAE, StyleGAN-XL and InsGen models trained on FFHQ samples.

### v2

@@ 6 Numerical Results
We evaluated the FKEA method on several image, text, and video datasets to assess its performance in quantifying diversity in different domains. In the experiments, we computed the empirical covariance matrix of $2r$-dimensional Fourier features with a Gaussian kernel with bandwidth parameter $\sigma$ tuned for each dataset, and then applied FKEA approximation for the $\mathrm{VENDI}_{1}$, $\mathrm{VENDI}_{1.5}$, and the $\mathrm{RKE}$ (same as $\mathrm{VENDI}_{2}$) scores. An algorithm to compute these scores is presented in Algorithm 1. Experiments were conducted on RTX3090 GPUs. We interpreted the modes identified by FKEA entropy-based diversity evaluation using the eigenvectors of the proxy covariance matrix as discussed in Remark 2. For each eigenvector, we presented the training data with maximum eigenfunction values corresponding to the eigenvector.
Time Complexity of FKEA metrics. To highlight the computational advantages of transitioning to FKEA, Table 1 presents a comparison of the metric computations for VENDI and RKE on the ImageNet dataset, with sample sizes ranging from 10k to 250k. Our results show that VENDI and RKE become computationally intractable due to memory overflow. In contrast, the FKEA method efficiently scales up to $n=250k$ samples, maintaining optimal computational time.
Table 1: Time complexity for FKEA and non-FKEA based metrics (RKE and VENDI) on ImageNet dataset with DinoV2 embedding. Computation of VENDI and RKE on 40k+ samples are omitted due to memory overflow during metric computation.
Time (sec)
$80002r=$ 8000 $160002r=$ 16000
Metric n=10k n=20k n=30k n=40k n=50k n=100k n=250k n=10k n=20k n=30k n=40k n=50k n=100k n=250k
FKEA-RKE 7 16 25 34 43 87 238 37 78 120 162 203 433 1138
FKEA-VENDI 11 19 27 37 45 104 267 48 89 130 173 213 459 1236
RKE 217 1324 4007 - - - - 218 1330 4021 - - - -
VENDI 286 1774 5488 - - - - 287 1780 5502 - - - -
@@
Experimental Results on Image Data. To investigate the FKEA method’s diversity evaluation in settings where we know the ground-truth clusters and their quantity, we simulated an experiment on the colored MNIST [46] data with the images of 10 colored digits as shown in Figure 2. We evaluated the FKEA approximations of the diversity scores while including samples from $t$ digits for $t\in\{1,\ldots,10\}$. The plots in Figure 2 indicate the increasing trend of the scores and FKEA’s tight approximations of the scores. Also, we show the top-20 training samples with the highest scores according to the top-10 FKEA eigenvectors, showing the method captured the ground-truth clusters.
@@
We conducted an experiment on the ImageNet data to monitor the scores’ behavior evaluated for 50k samples from an increasing number of ImageNet labels. Figure 3 shows the increasing trend of the scores as well as the top-9 samples representing the top-4 identified clusters used for the entropy calculation. Also, Figure 4 presents the FKEA approximated entropy scores with different truncation factors in StyleGAN3 [47] on 30k generated data for each truncation factor, showing the increasing diversity scores with the truncation factor. We defer discussing the results on AFHQ [48], MS-COCO [49], F-MNIST [50] datasets to the Appendix.
@@
Figure 3: RFF-based identified clusters used in FKEA Evaluation in ImageNet dataset with DinoV2 embedding, Fourier feature dimension $2r=16k$ and Gaussian Kernel bandwidth $\sigma=25$. The graphs indicate increase in FKEA diversity metrics with increasing number of labels per 50k samples.
@@
Experimental Results on Text and Video Data. To perform experiments on the text data with known clustering ground-truth, we generated 500,000 paragraphs using GPT-3.5 [51] about 100 randomly selected countries (5k samples per country). In the experiments, the text embedding used was text-embedding-3-large [52, 53, 51]. We evaluated the diversity scores over data subsets of size 50k with different numbers of mentioned countries. Figure 5 shows the growing trend of the diversity scores when including more countries. The figure also shows the countries mentioned in the top-6 modes provided by FKEA-based principal eigenvectors, which shows the RFF-based clustering of countries correlates with their continent and geographical location. We discuss the numerical results on non-synthetic text datasets, Wikipedia, CNN/Dailymail [54][55], CMU Movie Corpus [56], in the Appendix.
For video data experiments, we considered two standard video datasets, UCF101 [57] and Kinetics-400 [58]. Following the video evaluation literature [59, 60], we used the I3D pre-trained model [61] as embedding, which maps a video sample to a 1024-dimensional vector. As shown in Figure 6, increasing the number of video classes of test samples led to an increase in the FKEA approximated diversity metrics. Also, while the samples identified for the first identified cluster look broad, the next modes seemed more specific. We discuss the results of the Kinetics dataset in the Appendix.
@@ 7 Conclusion
Acknowledgments
The work of Farzan Farnia is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by a CUHK Direct Research Grant with CUHK Project No. 4055164. Xuenan Cao’s work is supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14602223. Andrej Bogdanov’s work is supported by an NSERC Discovery Grant.
@@ A.1 Proof of Theorem 1
The proof of Theorem 1 combines third ingredients. The first is the relation between the circuit size of a function $C$ and of its partial derivatives $\nabla C=(\partial C/\partial x_{1},\dots,\partial C/\partial x_{n})$.
@@
Let $C$ be a circuit over basis $B$ and $t$ be an input to $C$. There exists a circuit that computes the derivative $\partial g/\partial t$ for every gate $g$ of $C$ over basis $\nabla B\cup\{+,\times\}$ whose size is within a constant factor of the size of $C$.
@@
We first illustrate the proof in the special case when $\|X\|$ is within the radius of convergence of $f$. Namely, $f(x)$ is represented by the absolutely convergent series $\sum f^{(k)}(0)x^{k}/k!$ for all $\lvert x\rvert\leq\rho$. Then $f(X)=\sum f^{(k)}(0)X^{k}/k!$ assuming $\|X\|\leq\rho$. By linearity (and using the fact that derivatives preserve radius of convergence) it is sufficient to show that
@@
Assume $\mathrm{Tr}\rho^{\alpha}$ (resp., $-\mathrm{Tr}\rho\log\rho$) has circuit size $s(d)$. By Lemma 1 and Lemma 3, $\nabla\mathrm{Tr}\rho^{\alpha}=\alpha\rho^{\alpha-1}$ (resp., $-\nabla\mathrm{Tr}\rho\log\rho=\log\rho-1/\ln 2$) has circuit size $O(s(d))$. For every symmetric matrix $X$ and sufficiently small $t$, the matrix $\rho=I+tX$ is positive semi-definite. By Lemma 2 the $\mathbb{R}^{d^{2}}$-valued function $\partial^{2}\rho/\partial t^{2}$ has circuit size $O(s^{d})$. The value of this function at $t=0$ is $\alpha(\alpha-1)(\alpha-2)X^{2}$ (resp., $X^{2}$), namely the square of the input matrix $X$ up to constant. Finally, computing the product $AB$ reduces to squaring the symmetric matrix
@@ A.2 Proof of Theorem 2
=$\stackrel{{\scriptstyle(a)}}{{=}}\,\int\widehat{\kappa_{\sigma}}(\boldsymbol{\omega})\exp\bigl{(}i\boldsymbol{\omega}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\mathrm{d}\boldsymbol{\omega}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\,\int\widehat{\kappa_{\sigma}}(\boldsymbol{\omega})\cos\bigl{(}\boldsymbol{\omega}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\mathrm{d}\boldsymbol{\omega}$
$\stackrel{{\scriptstyle}}{{=}}\,\mathbb{E}_{\boldsymbol{\omega}\sim\widehat{\kappa}}\Bigl{[}\cos\bigl{(}\boldsymbol{\omega}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\Bigr{]}$
$=\,\mathbb{E}_{\boldsymbol{\omega}\sim\widehat{\kappa}}\Bigl{[}\cos\bigl{(}\boldsymbol{\omega}^{\top}\mathbf{x})\cos\bigl{(}\boldsymbol{\omega}^{\top}\mathbf{x}^{\prime})+\sin\bigl{(}\boldsymbol{\omega}^{\top}\mathbf{x})\sin\bigl{(}\boldsymbol{\omega}^{\top}\mathbf{x}^{\prime})\Bigr{]}$
@@
Therefore, since $\bigl{|}\cos\bigl{(}\boldsymbol{\omega}^{\top}\mathbf{y}\bigr{)}\bigr{|}\leq 1$ for all $\boldsymbol{\omega}$ and $\mathbf{y}$, one can apply Hoeffding’s inequality to show for independently drawn $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{r}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\widehat{\kappa}$ the following probably correct bound holds:
$\mathbb{P}\biggl{(}\Bigl{|}\frac{1}{r}\sum_{i=1}^{r}\cos\bigl{(}\boldsymbol{\omega}_{i}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}-\mathbb{E}_{\boldsymbol{\omega}\sim\widehat{\kappa}}\Bigl{[}\cos\bigl{(}\boldsymbol{\omega}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\Bigr{]}\Bigr{|}\geq\epsilon\biggr{)}\leq 2\exp\Bigl{(}-\frac{r\epsilon^{2}}{2}\Bigr{)}$
Therefore, as the identity $\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)$ reveals $\frac{1}{r}\sum_{i=1}^{r}\cos\bigl{(}\boldsymbol{\omega}_{i}^{\top}(\mathbf{x}-\mathbf{x}^{\prime})\bigr{)}\,=\,\widetilde{\phi}_{r}(\mathbf{x})^{\top}\widetilde{\phi}_{r}(\mathbf{x}^{\prime})$, the above bound can be rewritten as
@@
$\mathbb{P}\biggl{(}\max_{1\leq i,j\leq n}\Bigl{(}\widetilde{\phi}_{r}(\mathbf{x}_{i})^{\top}\widetilde{\phi}_{r}(\mathbf{x}_{j})-k_{\text{\rm Gaussian}(\sigma^{2})}(\mathbf{x}_{i},\mathbf{x}_{j})\Bigr{)}^{2}\geq\epsilon^{2}\biggr{)}\leq 2\binom{n}{2}\exp\Bigl{(}-\frac{r\epsilon^{2}}{2}\Bigr{)}.$
@@
$\mathbb{P}\Bigl{(}\bigl{\|}\frac{1}{n}\widetilde{K}-\frac{1}{n}K\bigr{\|}^{2}_{F}\,\geq\,n^{2}\frac{\epsilon^{2}}{n^{2}}\Bigr{)}\;\leq\;\binom{n}{2}\exp\Bigl{(}-\frac{r\epsilon^{2}}{2}\Bigr{)}.$
@@
Figure 7: RFF-based identified clusters used in FKEA Evaluation in ImageNet dataset with DinoV2 embeddings and bandwidth $\sigma=25$ at varying number of samples $n$
Appendix B Limitations
Appendix C Additional Numerical Results
C.1 Real Image Dataset Modes
This section details the results of cluster analyses conducted on various real-world datasets, including FFHQ, AFHQ, MSCOCO, and Fashion-MNIST. Each dataset’s results are organized into clusters identified by RFF method in FKEA evaluation.
Figure 8: RFF-based identified clusters used in FKEA Evaluation in FFHQ dataset with DinoV2 embeddings and bandwidth $\sigma=20$
Figure 9: RFF-based identified clusters used in FKEA Evaluation in AFHQ dataset with DinoV2 embeddings and bandwidth $\sigma=20$
Figure 10: RFF-based identified clusters used in FKEA Evaluation in Microsoft COCO dataset with DinoV2 embeddings and bandwidth $\sigma=22$
Figure 11: RFF-based identified clusters used in FKEA Evaluation in FASHION-MNIST [50] dataset with pixel embeddings and bandwidth $\sigma=15$
Figure 12: RFF-based identified clusters used in FKEA Evaluation in colored FASHION-MNIST [50] dataset with pixel embeddings and bandwidth $\sigma=4.5$
C.2 The effect of number of datapoints on clustering results with FKEA
In this section, we evaluate the quality of clusters obtained from the ImageNet dataset as the number of samples $n$ varies. Specifically, we compare clustering results for 10k, 50k, 100k, and 250k samples. Figure 7 illustrates the first four modes derived from the FKEA framework.
At $n=10k$, the clusters exhibit noise and often merge unrelated modes, as seen in Mode 2, where elephants and foxes appear within the same cluster. As $n$ increases, the clustering quality improves, becoming more coherent and meaningful. This trend is particularly evident in Modes 1 and 2, where the clusters more accurately reflect distinct semantic groups.
These findings highlight the importance of scaling VENDI and RKE scores, as computational overhead becomes a critical factor in assessing the diversity of generative models. Scaling these metrics allows for a more efficient evaluation, especially when dealing with large datasets and high sample counts.
C.3 Comparison between Generative Models on ImageNet dataset
In this section we report the FKEA scores for various generative models on ImageNet dataset. Table 3 evaluates the diversity scores of various ImageNet GAN models using the FKEA method applied to VENDI-1 and RKE, with potential extension to the entire VENDI family. The comparison includes baseline diversity metrics such as Inception Score [12], FID [7], Improved Precision/Recall [10], and Density/Coverage [11].
Dataset (100k) - - - - - - 9176.9 996.7
ADM [65] 542.6 11.12 0.78 0.79 0.88 0.89 8360.3 633.4
ADMG [65] 659.3 5.63 0.87 0.84 0.80 0.85 8524.2 811.5
ADMG-ADMU [65] 701.6 4.78 0.90 0.73 1.20 0.96 8577.6 839.8
BigGAN [66] 696.4 7.91 0.81 0.44 0.99 0.57 7120.5 492.4
DiT-XL-2 [67] 743.2 3.56 0.92 0.84 1.16 0.97 8626.5 855.8
GigaGAN [68] 678.8 4.29 0.89 0.74 0.74 0.70 8432.5 671.6
LDM [69] 734.4 4.75 0.93 0.76 1.04 0.93 8573.7 811.9
Mask-GIT [70] 717.4 5.66 0.91 0.72 1.01 0.82 8557.4 759.5
RQ-Transformer [71] 558.3 9.57 0.80 0.76 0.77 0.59 8078.4 512.1
StyleGAN-XL[72] 675.4 4.34 0.89 0.74 1.18 0.96 8171.9 703.5
C.4 Synthetic Image Dataset Modes
In addition to running clustering on ImageNet dataset, we also applied FKEA with varying Gaussian Kernel bandwidth parameter $\sigma$ to other datasets. The results are presented for FFHQ (Figure 8), AFHQ (Figure 9) Microsoft COCO (Figure 10) and Mono/Color versions of F-MNIST[50] (Figures 11 and 12) up to top 8 modes.
The experimental setup is similar to figure 3 with the only change is optimised bandwidth for each dataset, since datasets differ in number and typicality of the samples.